### 26 Jan 2024 v1

@@ 7 Summary
In this paper, we analyse the impact of the environment on the evolution of (un)barred disc galaxies at $z=1$. We use the TNG50 simulation, which is a magneto-cosmological hydrodynamical simulation based on the $\Lambda$CDM model as part of the IllustrisTNG project (Pillepich et al. 2019; Nelson et al. 2019b).
The galaxy components such as the classical bulge, thin disc, and thick disc are identified using the kinematic decomposition software mordor (Zana et al. 2022) and the identification of the bar structures is done using the Fourier decomposition of the face-on stellar surface density.
To quantify the environment, we investigate the merger histories of disc galaxy descendants and identify the temporal region of influence of major mergers in the galaxy. We also explore the distance to the closest massive companion to identify the possible tidal interactions that galaxies experience, and, finally, we investigate the number of satellites as a function of redshift to assess the large-scale environment.
Our findings include the following:
@@
  • Those galaxies that transform into lenticulars and spheroids have an active (major) merger history (see Fig. 4 and Table 2). The descendant spheroidals and lenticulars exhibit the highest fraction of galaxies that experience at least one major merger ($0.36$ $0.38$), while the descendant disc galaxies have the lowest fraction ($0.15$) since $z=1$. Furthermore, spheroidal galaxies generally exist in a more dense environment compared to other galaxies, as they possess the highest fraction of satellite galaxies over time.
To explore the rise and decline of bar structures, we divide the parent disc sample into two categories: barred disc galaxies at $z=1$ and unbarred disc galaxies at the same redshift. Subsequently, we monitor the development of these two groups. Concerning the barred disc galaxies at $z=1$, our findings indicate that:
  • $5.4$ and $1.1$ per cent of disc and lenticular descendants, respectively, lose their bar (see Table 1 and Fig. 18). On the other hand, $33.3$ and $26.9$ per cent of the barred disc and lenticular descendants maintain their bars, respectively. The remaining $33.3$ per cent become spheroidals.
  • There are variations in the characteristics of the bars: the galaxies that are capable of retaining their bars exhibit the strongest and elongated bars, whereas the galaxies whose bar dissolves overtime overall tend to have weaker and shorter bars to start with at $z=1$ (see Figs. 6 & 7). The subsample that preserves its bar structure has its stellar disc dominated by the thin disc component.
@@
  • $53.1$ per cent of unbarred descendant galaxies develop a bar (Table 1 and Fig. 18), whereas $27.1$ of them do not develop a stable bar at $z=0$ and $19.8$ per cent transform into a spheroidal.
@@
  • In disc galaxies that experience bar formation, the bar structure typically emerges within the first Gyrs after $z=1$ (see Fig. 12). These bars are relatively weaker and shorter compared to the ones of disc galaxies that have a bar formed at earlier times. This observation aligns with the notion that the bars in the former galaxies formed at an early stage (Rosas-Guevara et al. 2020; Izquierdo-Villalba et al. 2022). The lenticular descendants form the bar before undergoing morphological transformation, and the disc has not been destroyed yet (see Fig. 13).
@@
  • Unbarred disc descendants that fail to create a stable bar exhibit the lowest galaxy fraction that has undergone major mergers ($0.24$, see Fig. 14). Conversely, unbarred disc descendants that eventually form a stable bar have the highest galaxy fraction ($0.36$) that has undergone major mergers, excluding those that transform into lenticulars and never undergo bar formation.
@@
  • Galaxies that form a bar after $z=0$ show a temporal correlation with a recent major merger.The results show that at least $30$ per cent and $20$ per cent of the disc and lenticular galaxies form their bars during the influence of the major merger (see Fig. 15). We found similar results when we looked at the time when the minor mergers happened.
@@
  • We find, on average, there is no difference in the distances to the closest massive companion between galaxies that form a bar and those that do not (see Fig. 16).
@@
  • On average, galaxies that form a bar structure exhibit a greater stellar mass (e.g. Gavazzi et al. 2015; Consolandi 2016), a reduced gas fraction (e.g. Masters et al. 2012; Kruk et al. 2018), and an elevated ratio of stellar mass to dark matter mass (e.g. Fragkoudi et al. 2021; Izquierdo-Villalba et al. 2022). Regarding the final one, there is a possible connection with the merger history, both before and after $z=1$. We find a disparity in the ratio of stellar mass to dark matter mass for a specific stellar mass. This disparity was already higher in barred galaxies, even before the formation of the bars. Additionally, unbarred galaxies that form a bar after $z=1$ undergo a higher number of major mergers compared to unbarred galaxies that never develop a bar. This suggests that the formation of barred galaxies occurs before the formation of galaxies that never generated a bar.
@@
In summary, we find that bars are prevalent in disc galaxies and are stable. Nevertheless, bar evolution can be influenced by the surrounding environment. Tidal interactions resulting from mergers or encounters with close companions have the potential to disrupt the central region of the disc. This disruption could manifest as either the dissolution of the bar, the formation of a bar, or, in the most severe case, the disc complete destruction, which would cause a morphological transformation of the galaxy. Weak and short bars at $z=1$ that enter crowded environments or experience a merger may dissolve, according to our findings. Unbarred galaxies at $z=1$ that undergo major/minor mergers may eventually develop a bar. The influence of a merger or close encounter on the formation or dissolution of a bar is dependent on the orbital parameters, which could be investigated in future studies through the combination of cosmological hydrodynamic simulations and high-temporal resolution simulations. In conclusion, our findings suggest that the environment may have influenced the evolution of the bar fraction, a commonly observable used to understand the evolution of bars.
@@ Acknowledgements.
The authors thank the TNG team for early access to the TNG50 data prior to its public release. YRG acknowledges the support of the“Juan de la Cierva Incorporation” Fellowship (IJC2019-041131-I). S.B. acknowledges support from the Spanish Ministerio de Ciencia e Innovación through project PID2021-124243NB-C21. D.I.V. acknowledges the financial support provided under the European Union’s H2020 ERC Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691).
@@ Appendix A Calculation of the dissolution time of a bar
In this section, we present some examples of disc galaxies that dissolved their bar and the time of bar dissolution. As explained in subsection 2.2. The time dissolution is determined by the lookback time when the bar strength is smaller than 0.2 and the bar length is smaller than the minimum radius imposed due to spatial resolution. Fig 19 shows two examples of the bar strength and bar length of disc galaxies that were barred at $z=1$ and dissolved their bar after.
Figure 19: Examples of the bar evolution of disc galaxies at $z=1$ that dissolved their bar. The black solid line indicates the time of the dissolution of the bar calculated, and the grey solid line represents the lookback time that corresponds to $z=1$.

### 07 Feb 2024 v2

@@ 7 Summary
In this study, we analysed the impact of the environment on the evolution of barred and unbarred disc galaxies at $z=1$. We used the TNG50 simulation, which is a magneto-cosmological hydrodynamical simulation based on the $\Lambda$CDM model as part of the IllustrisTNG project (Pillepich et al. 2019; Nelson et al. 2019b).
As part of the approach taken here, the galaxy components, such as the classical bulge, thin disc, and thick disc, are identified using the kinematic decomposition software mordor (Zana et al. 2022) and the identification of the bar structures is done using the Fourier decomposition of the face-on stellar surface density.
To quantify the environment, we investigated the merger histories of disc galaxy descendants and identified the temporal region of influence of major mergers in the galaxy. We also explored the distance to the closest massive companion to identify the possible tidal interactions that galaxies experience, and, finally, we investigated the number of satellites as a function of redshift to assess the large-scale environment.
Our findings can be summarised as follows:
@@
  • Those galaxies that transform into lenticular and spheroidal galaxies have an active (major) merger history (see Fig. 4 and Table 2). The descendant spheroidal and lenticular galaxies exhibit the highest fraction of galaxies that experience at least one major merger ($0.36$ and $0.38$), while the descendant disc galaxies have the lowest fraction ($0.15$) since $z=1$. Furthermore, spheroidal galaxies generally exist in a denser environment compared to other galaxies, as they possess the highest fraction of satellite galaxies over time.
To explore the rise and decline of bar structures, we divided the parent disc sample into two categories: barred disc galaxies at $z=1$ and unbarred disc galaxies at the same redshift. Subsequently, we monitored the development of these two groups. Concerning the barred disc galaxies at $z=1$, our findings indicate that:
  • $5.4$ and $1.1$ per cent of disc and lenticular descendants, respectively, lose their bar (see Table 1 and Fig. 18). On the other hand, $33.3$ and $26.9$ per cent of the barred disc and lenticular descendants, respectively, maintain their bars. The remaining $33.3$ per cent become spheroidal galaxies.
  • There are variations in the characteristics of the bars: the galaxies that are capable of retaining their bars exhibit the strongest and most elongated bars, whereas the galaxies whose bar dissolves overtime overall tend to have weaker and shorter bars to start with at $z=1$ (see Figs. 6 and 7). The subsample that preserves its bar structure has its stellar disc dominated by the thin-disc component.
@@
  • $53.1$ per cent of unbarred descendant galaxies develop a bar (Table 1 and Fig. 18), whereas $27.1$ of them do not develop a stable bar at $z=0$ and $19.8$ per cent transform into a spheroidal galaxy.
@@
  • In disc galaxies that experience bar formation, the bar structure typically emerges within the first few gigayears after $z=1$ (see Fig. 12). These bars are relatively weak and short compared to the bars of disc galaxies that formed a bar at earlier times. This observation aligns with the notion that the bars in the former galaxies formed at an earlier stage (Rosas-Guevara et al. 2020; Izquierdo-Villalba et al. 2022). The lenticular descendants form the bar before undergoing morphological transformation, and the disc has not yet been destroyed (see Fig. 13).
@@
  • Unbarred disc descendants that fail to create a stable bar exhibit the lowest galaxy fraction that has undergone major mergers ($0.24$, see Fig. 14). Conversely, unbarred disc descendants that eventually form a stable bar have the highest galaxy fraction ($0.36$) that has undergone major mergers, excluding those that transform into lenticular galaxies and never undergo bar formation.
@@
  • Galaxies that form a bar after $z=0$ show a temporal correlation with a recent major merger. Our results show that at least $30$ per cent and $20$ per cent of the disc and lenticular galaxies, respectively, form their bars during the influence of the major merger (see Fig. 15). We find similar results when we look at the time when the minor mergers happened.
@@
  • We find that, on average, there is no difference in the distances to the closest massive companion between galaxies that form a bar and those that do not (see Fig. 16).
@@
  • On average, galaxies that form a bar structure exhibit a greater stellar mass (e.g. Gavazzi et al. 2015; Consolandi 2016), a reduced gas fraction (e.g. Masters et al. 2012; Kruk et al. 2018), and an elevated ratio of stellar mass to dark matter mass (e.g. Fragkoudi et al. 2021; Izquierdo-Villalba et al. 2022). There is a possible connection between the evolution of this latter ratio and the merger history both before and after $z=1$. We find a disparity in the ratio of stellar mass to dark matter mass for a specific stellar mass. This disparity is already higher in barred galaxies, even before the formation of the bars. Additionally, unbarred galaxies that form a bar after $z=1$ undergo a higher number of major mergers compared to unbarred galaxies that never develop a bar. This suggests that the formation of barred galaxies occurs before the formation of galaxies that never generate a bar.
@@
In summary, we find that bars are prevalent in disc galaxies and are stable. Nevertheless, bar evolution can be influenced by the surrounding environment. Tidal interactions resulting from mergers or encounters with close companions have the potential to disrupt the central region of the disc. This disruption can manifest as either the dissolution of the bar, the formation of a bar, or, in the most severe case, the complete destruction of the disc, which causes a morphological transformation of the galaxy. Weak and short bars at $z=1$ that enter crowded environments or experience a merger may dissolve according to our findings. Unbarred galaxies at $z=1$ that undergo major or minor mergers may eventually develop a bar. The influence of a merger or close encounter on the formation or dissolution of a bar is dependent on the orbital parameters, which could be investigated in future studies through the combination of cosmological hydrodynamic simulations and high-temporal-resolution simulations. In conclusion, our findings suggest that the environment may have influenced the evolution of the bar fraction, a common observable used to understand the evolution of bars.
@@ Acknowledgements.
The authors thank the referee for the useful comments and the TNG team for early access to the TNG50 data prior to its public release. Y.R.G. acknowledges the support of the“Juan de la Cierva Incorporation” Fellowship (IJC2019-041131-I). Y.R.G. thanks Karin Menendez-Delmestre and the Conference for Galactic Bars: Driving and Decoding Galaxy Evolution for the useful discussions. S.B. acknowledges support from the Spanish Ministerio de Ciencia e Innovación through project PID2021-124243NB-C21. D.I.V. acknowledges the financial support provided under the European Union’s H2020 ERC Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691).
@@ Appendix A Calculation of the dissolution time of a bar
In this section, we present some examples of disc galaxies that dissolved their bar and the time of bar dissolution, as explained in subsection 2.2. The time dissolution is determined by the lookback time when the bar strength is smaller than 0.2 and the bar length is smaller than the minimum radius imposed due to spatial resolution. Figure 19 shows two examples of the bar strength and bar length of disc galaxies that were barred at $z=1$ and dissolved their bar thereafter.
Figure 19: Examples of the bar evolution of disc galaxies at $z=1$ that dissolved their bar. The black solid line indicates the calculated time of the dissolution of the bar, and the grey solid line represents the lookback time that corresponds to $z=1$.